OpenStudy (shiree19):

Please HELP!! Fan & Medal! How do you graph y = 1/2 * 11 – 6? I cannot use a graphing calculator. Thank you!

OpenStudy (shiree19):

how do you graph y = 1/2 * 11 – 6

OpenStudy (shiree19):

Errrr it won't work.\[y = \frac{ 1 }{ 2 } \times 11 - 6\]